# Autonomous Constellation Fault Monitoring with Inter-satellite Links: A Rigidity-Based Approach

### v1

@@ Abstract
To address the need for robust positioning, navigation, and timing services in lunar and Martian environments, this paper proposes a novel fault detection framework for satellite constellations using inter-satellite ranging (ISR). Traditional fault monitoring methods rely on intense monitoring from ground-based stations, which are impractical for lunar and Martian missions due to cost constraints. Our approach leverages graph-rigidity theory to detect faults without relying on precise ephemeris. We model satellite constellations as graphs where satellites are vertices and inter-satellite links are edges. By analyzing the Euclidean Distance Matrix (EDM) derived from ISR measurements, we identify faults through the singular values of the geometric-centered EDM (GCEDM). A neural network predictor is employed to handle the diverse geometry of the graph, enhancing fault detection robustness. The proposed method is validated through simulations of constellations around Mars and the Moon, demonstrating its effectiveness in various configurations. This research contributes to the reliable operation of satellite constellations for future lunar and Martian exploration missions.
@@ 1 Introduction
To meet the growing need for robust positioning, navigation, and timing (PNT) services at the lunar surface and lunar orbits, NASA and its international partners are collaborating to develop LunaNet (Israel et al.,, 2020), a network of networks providing data relay, PNT, detection, and science services. In LunaNet, these services are provided through cooperation among interoperable systems that evolve over time to meet the growing needs for these services, efficiently establishing a reliable, sustainable, and scalable network. NASA also envisions extending LunaNet to MarsNet in the future (National Aeronautics and Space Administration,, 2023).
@@
It is crucial to monitor LunaNet navigation satellites for the reliable operation of safety-critical missions. The quality of lunar navigation signals can be compromised by various system faults, such as clock runoffs (e.g. phase and frequency jumps (Weiss et al.,, 2010)), unflagged maneuvers, failures in satellite payload signal generation components, and code-carrier incoherence (Working Group C, ARAIM Technical Subgroup,, 2012). For terrestrial GNSS, the system fault alerts are provided as integrity information from GNSS augmentation systems, such as the satellite-based augmentation system (SBAS) (Van Diggelen,, 2009). There also exist algorithms to monitor and remove faults on the receiver side, known as receiver autonomous integrity monitoring (RAIM) (MISRA et al.,, 1993). However, given the limited resources of lunar missions, it is preferable to perform fault monitoring within the navigation system. While the LunaNet Relay Service Documents (SRD) state that robustness of the signal should be a key consideration for LunaNet (National Aeronautics and Space Administration,, 2022), the specific methodology to monitor faults on LunaNet satellites is yet solidified.
@@
Satellite fault monitoring in LunaNet (or MarsNet) can be challenging because there are no dedicated monitoring station networks on the Moon (or Mars), especially in the early stage of operation. On Earth, SBAS monitors satellite faults by collecting GNSS signals at monitoring stations that are accurately located, and processes them in a central computing center where differential correction and integrity messages are calculated (Van Diggelen,, 2009). However, navigation satellite systems on the Moon or Mars will likely have none or very few monitoring stations on the planet’s surface to monitor the signals, due to stringent cost constraints to deploy and maintain these stations. Therefore, onboard algorithms that can monitor satellite faults onboard are desired for LunaNet and MarsNet.
@@
One promising approach for autonomous satellite fault monitoring is to use inter-satellite ranging (ISR), a concept that has been investigated for terrestrial GNSS constellations (Wolf,, 2000; Rodríguez-Pérez et al.,, 2011). The proposed autonomous satellite fault monitoring algorithms from previous works assume the availability of precise ephemeris to compute the estimated range between the satellites, which is used to compute the residuals that are over-bounded by Gaussian distributions. However, obtaining a precise ephemeris is challenging for future lunar and Martian constellations due to the limited number of monitoring stations and the lower stability of onboard clocks due to stringent cost requirements. Moreover, if the ISR measurements are used for orbit determination and time synchronization (ODTS) to generate the ephemeris of the navigation satellites as proposed in the LunaNet concept, this would create a chicken-and-egg problem of ODTS and satellite fault detection. The previously proposed algorithms also assume that the ISR measurements are sufficiently precise, well-calibrated, and contain no faults, which may be too strong of an assumption in practice.
@@
Our algorithm uses the ISR bias generated from faults to detect a deformation in the 3D rigid graph. We start by modeling satellite constellations with ISR measurements as a graph, where satellites are modeled as vertices and links between satellites are modeled as edges with weights as their measured ranges. When the constellation size is sufficiently large and a meshed network within the constellation can be constructed, this graph contains multiple subgraphs of $k$ satellites that are fully connected, which are called $k$-cliques. $k$-cliques of $k\geq 5$ are known to be 2-vertex rigid in 3-dimensional space, which means they remain rigid (graphs cannot be susceptible to continuous flexing are called rigid graphs), after removing any vertex from the graph (Alireza Motevallian et al.,, 2015). Therefore, by checking the consistency of the range measurements for each of the $k$-clique subgraphs with $k\geq 5$, we can determine if there is a fault satellite within the nodes that adds biases to its connected edges, which makes the graph un-realizable in 3-dimensional space.
@@
In particular, our proposed method monitors the singular values of the geometric-centered Euclidean Distance Matrix (GCEDM) (Dokmanić et al.,, 2015) constructed from the range measurements. This is inspired by the GNSS fault detection algorithm by (Knowles and Gao,, 2024), where they propose to monitor the 4th and 5th singular value of the GCEDM, constructed from the observed range between the user and the GNSS satellite and the ephemeris of the GNSS satellites to identify GNSS satellites with faults. Using the fact that the 4th and 5th singular values increase when a fault is present, their method identifies the presence of a fault satellite whenever the sum of these two singular values goes over a certain threshold. This method is promising for fault detection, because it does not require solving for the user position to identify faults, and resolves the aforementioned chicken-and-egg problem. However, in our setting fault detection is more challenging since we have noisy ranges between all nodes, while in (Knowles and Gao,, 2024), it is assumed that the ranges between GNSS satellites are known very precisely. In addition, the relative geometry between the nodes is more diverse in our setting, since the inter-satellite signals can come from all directions, compared to GPS satellite signals which only come over the horizon in GPS user cases.
To overcome this challenge, we propose a fault detection method based on a neural network predictor that predicts the tail distribution of the singular values of the GCEDM from the first three singular values and their corresponding singular vectors. This enables the fault detection algorithm to be applied to fully connected subgraphs in dynamic scenarios where fixed thresholds do not work well
@@
The contribution of this paper is summarized below.
@@
  • We prove several key properties about the ranks of EDMs and GCEDMs to provide mathematical backing of fault detection using the fourth and fifth singular values of the GCEDM.
  • We propose a novel machine-learning approach to predict tails of the distribution of the singular values of the noisy GCEDMs, that can be used for fault satellite detection.
  • We propose a fault detection framework that combines $k$-clique finding with the neural network predictor, that can be applied to dynamically changing graph topologies and relative geometries.
@@
  • We validate our algorithm in a simulated constellation around Mars and the Moon. We show how the hyper-parameters and fault magnitudes affect the fault detection performance.
@@
The paper is arranged as follows. Section 2 introduces the notations and proves some properties of the rank of the EDMs and GCEDMs. In section 3, the fault detection framework is proposed. In section 4, performance evaluation for Mars and Lunar constellations are provided. The paper concludes in section 5.
@@ 2 Properties of the Euclidean Distance Matrix
In this section, we introduce and prove several properties of the EDM and GCEDM constructed from the observed ranges between the satellites.
2.1 Definitions
Consider a fully connected graph of $n$ nodes, which corresponds to satellites. Let $\bm{x}_{i}\in\mathbb{R}^{d},(i\in\{1,\ldots,n\})$ be the position of the $i$th node, and $\mathbf{X}\in\mathbb{R}^{d\times n}$ be a matrix collecting these points. In this paper, we are interested in the 3D case, $d=3$. By establishing two-way inter-satellite links, we obtain the range measurement $r_{ij}$ between the two satellites $i$ and $j$,
$r_{ij}=\begin{cases}\|\bm{x}_{i}-\bm{x}_{j}\|+w_{ij}+b_{ij}&(i\neq j)\\ 0&(i=j)\end{cases}$ (1)
Above, $w_{ij}$ is the range measurement noise, and $b_{ij}$ is the bias of the range measurements. We assume the measurement noise $w_{ij}$ is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma_{w})$. The bias term is determined as follows
$b_{ij}$ $=b_{i}+b_{j},$ (2)
$b_{k}$ =$=\begin{cases}\bar{b}&\text{satellite k is fault}\\ 0&\text{satellite k is normal}\end{cases},\quad k\in\{i,j\}$ (3)
[…]
Using the set of observed ranges $r_{ij}$, we construct an EDM, where its elements are equivalent to the square of the observed ranges ($r_{ij}^{2})$. Let $\mathbf{D}^{n,d,m}\in\mathbb{R}^{n\times n}$ be an EDM without measurement noise ($w_{ij}=0$) and $m$ nodes out of $n$ nodes being fault. For example, a noiseless EDM with 6 satellites (in 3D space) with 2 fault satellites is denoted as $\mathbf{D}^{6,3,2}$. Similarly, we define $\tilde{\mathbf{D}}^{n,d,m}\in\mathbb{R}^{n\times n}$ as EDM with measurement noise and $m$ nodes out of $n$ nodes having a fault. A GCEDM $\mathbf{G}^{n,d,m}$ (or $\tilde{\mathbf{G}}^{n,d,m}$ for noisy EDM $\tilde{\mathbf{D}}^{n,d,m}$) is constructed from the EDM from the following operation

### v2

@@ Abstract
To address the need for robust positioning, navigation, and timing services in lunar environments, this paper proposes a novel fault detection framework for satellite constellations using inter-satellite ranging (ISR). Traditionally, navigation satellites can depend on a robust network of ground-based stations for fault monitoring. However, due to cost constraints, a comprehensive ground segment on the lunar surface is impractical for lunar constellations. Our approach leverages vertex redundantly rigid graphs to detect faults without relying on precise ephemeris. We model satellite constellations as graphs where satellites are vertices and inter-satellite links are edges. We identify faults through the singular values of the geometric-centered Euclidean distance matrix (GCEDM) of 2-vertex redundantly rigid sub-graphs. The proposed method is validated through simulations of constellations around the Moon, demonstrating its effectiveness in various configurations. This research contributes to the reliable operation of satellite constellations for future lunar exploration missions.
@@ 1 Introduction
To meet the growing need for robust positioning, navigation, and timing (PNT) services at the lunar surface and lunar orbits, NASA and its international partners are collaborating to develop LunaNet (Israel et al.,, 2020), a network of networks providing data relay, PNT, detection, and science services. In LunaNet, these services are provided through cooperation among interoperable systems that evolve over time to meet the growing needs for these services, efficiently establishing a reliable, sustainable, and scalable network.
@@
It is crucial to monitor LunaNet navigation satellites for the reliable operation of safety-critical missions. The quality of lunar navigation signals can be compromised by various system faults, such as clock runoffs (e.g. phase and frequency jumps (Weiss et al.,, 2010)), unflagged maneuvers, failures in satellite payload signal generation components, and code-carrier incoherence (Walter et al.,, 2018). For terrestrial GNSS, the system fault alerts are provided as integrity information from GNSS augmentation systems, such as the satellite-based augmentation system (SBAS) (Van Diggelen,, 2009). There also exist algorithms to monitor and remove faults on the receiver side, known as receiver autonomous integrity monitoring (RAIM) (MISRA et al.,, 1993). However, given the limited resources of lunar missions, it is preferable to perform fault monitoring within the navigation system. While the LunaNet Relay Service Documents (SRD) state that robustness of the signal should be a key consideration for LunaNet (National Aeronautics and Space Administration,, 2022), the specific methodology to monitor faults on LunaNet satellites is yet solidified.
@@
Satellite fault monitoring in LunaNet can be challenging because there are no dedicated monitoring station networks on the Moon, especially in the early stage of operation. On Earth, SBAS monitors satellite faults by collecting GNSS signals at monitoring stations that are accurately located, and processes them in a central computing center where differential correction and integrity messages are calculated (Van Diggelen,, 2009). However, navigation satellite systems on the Moon or Mars will likely have none or very few monitoring stations on the planet’s surface to monitor the signals, due to stringent cost constraints to deploy and maintain these stations. Therefore, onboard algorithms that can monitor satellite faults are desired for LunaNet.
@@
One promising approach for autonomous satellite fault monitoring is to use inter-satellite ranging (ISR), a concept that has been investigated for terrestrial GNSS constellations (Wolf,, 2000; Rodríguez-Pérez et al.,, 2011). The proposed autonomous satellite fault monitoring algorithms from previous works assume the availability of precise ephemeris to compute the estimated range between the satellites, which is used to compute the residuals over-bounded by Gaussian distributions. However, obtaining a precise ephemeris is challenging for future lunar and Martian constellations due to the limited number of monitoring stations and the lower stability of onboard clocks due to stringent cost requirements. Moreover, if the ISR measurements are used for orbit determination and time synchronization (ODTS) to generate the ephemeris of the navigation satellites as proposed in the LunaNet concept, this would create a chicken-and-egg problem of ODTS and satellite fault detection. The previously proposed algorithms also assume that the ISR measurements are sufficiently precise, well-calibrated, and contain no faults, which may be too strong of an assumption in practice.
@@
Our algorithm uses the ISR bias generated from faults to detect a deformation in the 3D rigid graph. We start by modeling satellite constellations with ISR measurements as a graph, where satellites are modeled as vertices and links between satellites are modeled as edges with weights as their measured ranges. When the constellation size is sufficiently large and a meshed network within the constellation can be constructed, this graph contains multiple subgraphs of $k$ satellites that are fully connected, which are called $k$-cliques. $k$-cliques of $k\geq 5$ are known to be 2-vertex rigid in 3-dimensional space, which means they remain rigid (graphs cannot be susceptible to continuous flexing are called rigid graphs), after removing any vertex from the graph (Alireza Motevallian et al.,, 2015). We show that if a graph is 2-vertex rigid in 3-dimensional space, we can detect a fault satellite by checking if the given set of ranges is realizable in 3-dimensional space. Therefore, we can determine if there is a fault satellite within the nodes by checking if the given set of ranges is realizable in 3-dimensional space, for each of the $k$-clique subgraphs with $k\geq 5$.
@@
Our proposed method monitors the singular values of the geometric-centered Euclidean Distance Matrix (GCEDM) (Dokmanić et al.,, 2015) constructed from the range measurements, to check if a given set of ranges are realizable in 3-dimensional space. This is inspired by the GNSS fault detection algorithm by (Knowles and Gao,, 2023), where they propose to monitor the 4th and 5th singular value of the GCEDM, constructed from the observed range between the user and the GNSS satellite and the ephemeris of the GNSS satellites to identify GNSS satellites with faults. Using the fact that the 4th and 5th singular values increase when a fault is present, their method identifies the presence of a fault satellite whenever the sum of these two singular values goes over a certain threshold. In this paper, we prove why the 4th and 5th singular value increases under the presence of fault, and analyze the distribution of the singular values under the presence of noise in the range measurement. Using GCEDM is promising for fault detection because it does not require solving for the user position to identify graph realizability (and faults) and resolves the aforementioned chicken-and-egg problem.
@@
The contribution of this paper is summarized below. This work is based on our prior conference paper presented at the 2024 ION GNSS+ (Iiyama et al.,, 2024).
  • We propose a rigidity-based online fault detection framework for satellite constellations that do not require precise ephemeris or observations at the monitoring stations.
  • We identify the required graph topology to identify fault satellites from a set of measured inter-satellite ranges. In particular, we show that the graph has to be 2-vertex redundantly rigid to detect fault satellites.
@@
  • We prove several key properties about the ranks of EDMs and GCEDMs to provide mathematical backing for fault detection using the fourth and fifth singular values of the GCEDM.
@@
  • We validate our algorithm in a simulated constellation around the Moon. We show how the hyper-parameters and fault magnitudes affect the fault detection performance.
@@
The paper is arranged as follows. Section 2 discuss the required graph topology to detect satellite faults from observed ranges. Section 3 introduces the notations and proves some properties of the rank of the EDMs and GCEDMs. In section 4, the fault detection framework is proposed. In section 5, performance evaluation for lunar constellations are provided. The paper concludes in section 6.
2 Required Graph Topology for Fault Detection
In this section, we identify the required graph topology for fault detection.
2.1 Definitions
2.1.1 Range Measurements and Satellite Faults
Consider a graph of $n$ nodes, which corresponds to satellites. Let $\bm{x}_{i}\in\mathbb{R}^{d},(i\in\{1,\ldots,n\})$ be the position of the $i$th node, and $\mathbf{X}\in\mathbb{R}^{d\times n}$ be a matrix collecting these points. In this paper, we are interested in the 3D case, $d=3$. By establishing two-way inter-satellite links, we obtain the range measurement $r_{ij}$ between the two satellites $i$ and $j$,
$r_{ij}=\begin{cases}\|\bm{x}_{i}-\bm{x}_{j}\|+w_{ij}+f_{ij}=\frac{c}{2}\tau_{ij}&(i\neq j)\\ 0&(i=j)\end{cases}$ (1)
Above, $w_{ij}$ is the range measurement noise, $f_{ij}$ is the bias of the range measurements, $c$ is the light speed, and $\tau_{ij}$ is the (measured) two-way travel time. The noise $w_{ij}$ includes thermal noise in the receiver, un-calibrated instrumental delay, and light-time delay estimation errors. We assume the measurement noise $w_{ij}$ is sampled from a Gaussian distribution $\mathcal{N}(0,\sigma_{w})$. The bias term is determined as follows
$f_{ij}$ $=f_{i}+f_{j},$ (2)
$f_{k}$ =$=\begin{cases}\bar{f}&\text{satellite k is fault}\\ 0&\text{satellite k is normal}\end{cases},\quad k\in\{i,j\}$ (3)
which means we assume that the bias is zero when no satellites are in fault status. In other words, we assume a constant bias $\bar{f}$ gets inserted if either of the satellites in the edges is faulty.
2.1.2 Weighed Graph and Embeddability
Based on the graph and observed ranges, we define a weighted graph $G=\langle V,E,W\rangle$. A realization (embedding) of $G$ is d-dimensional Euclidean distance space, $\mathbb{R}^{d}$, is defined as a function $f_{r}$ that maps $V$ into $\mathbb{R}^{d}$, such that for each edge $e=(v_{i},v_{j})\in E$, $W(e)=r_{ij}$ (Jian et al.,, 2010).
A weighted graph $G=\langle V,E,W\rangle$ is called d-embeddable when there is a realization (embedding) $f:V\rightarrow\mathbb{R}^{d}$ that maps the edges to points ($\bm{x}_{1},\ldots,\bm{x}_{n}$) in $\mathbb{R}^{d}$ space.
A weighted graph is fault disprovable if and only if the embeddability of $G$ implies that it contains no fault satellites.
2.1.3 Rigid and vertex redundantly rigid graphs
A graph is called rigid if it has no continuous deformation other than rotation, translation, and reflection while preserving the relative distance constraints between the vertices (Laman,, 1970). Otherwise, it is called a flexible graph. A graph is called k-vertex redundantly rigid if it remains rigid after removing any $(k-1)$ vertices. Similarly, if it remains rigid after removing any $(k-1)$ edges, it is called k-edge redundantly rigid.
(a) Flexible Graph
(b) Rigid Graph
(c) 2-vertex redundantly rigid graph
Figure 1: Flexible, Rigid, and 2-vertex redundantly rigid graphs in $\mathbb{R}^{2}$
2.2 Embeddability of Graphs with Fault Satellites
In this section, we prove the sufficient and necessary condition for the graph to be fault disprovable in $\mathbb{R}^{d}$. In the following proofs, we assume that there is no noise $w_{ij}$ in the range measurements, and $\bar{f}$ is an arbitrary continuous random variable.
2.2.1 Sufficient condition for fault disprovability
Proposition 2.1.
Suppose there is a single fault satellite on a graph $G^{\prime}=\langle V,E,W\rangle$ that is k-vertex ($k\geq 2$) redundantly rigid in $\mathbb{R}^{d}$. If $G$ is $d$-embeddable, then $G$ contains no fault vertex (satellite) with probability 1.
Proof.
Assume $G$ is k-vertex ($k\geq 2$) redundantly rigid, $d$-embeddable, and contains a fault satellite. Let $v$ a fault satellite in $G$ and $u\neq v$ be another (normal) satellite in $G$. Because $G$ is $k$-vertex ($k\geq 2$) redundantly rigid, after removing $v$, it is still rigid, which means there are finite discontinuous realizations up to congruence for $G^{v}=G\setminus{v}$. Similarly, there are finite discontinuous realizations up to congruence for $G^{u}=G\setminus{u}$. In each realization of $G^{v}$ and $G^{u}$, the distance between node $v$ and node $u$ is fixed. So, in total, there are finite discrete values, denoted by set $S$, for the distance between $u$ and $v$. As $G$ is embeddable and $e_{uv}=(u,v)\in E$, the measured distance of $e_{uv}$ is in $S$. However, since $e_{uv}$ is an edge connected to a fault satellite, the measured distance of $e_{uv}$ is a continuous random variable, so with a probability of 0, $W(e)\in S$, since $S$ has measure 0. So with a probability of 1, there is no fault satellite in $G$. ∎
The visual description of the proof is shown in Figure 2.
Figure 2: A visual description of the proof of Proposition 2.1 for $d=2$.
Proposition 2.2.
Given a weighted graph $G$, if $G$ is $k$-vetex redundantly rigid ($k\geq 2)$, then $G$ is fault disprovable.
Proof.
We can also prove that for arbitrary graphs, if $G$ is not embeddable, then $G$ contains a fault satellite. Therefore, with 2.1, if $G$ is $k$-vetex ($k\geq 2)$ redundantly rigid, then $G$ is fault disprovable. ∎
2.2.2 Necessary condition for fault disprovability
Next, we prove the necessary condition.
Proposition 2.3.
Given a weighted graph $G$, if $G$ is fault disprovable, then $G$ is $k$-edge redundantly rigid ($k\geq 2)$.
Proof.
We will prove the contraposition of the proposition: If $G$ is not $k$-edge redundantly rigid ($k\geq 2)$, $G$ is not fault disprovable.
Assume $G$ is not $k$-edge redundantly rigid and contains fault satellite. Then it is also not $k$-vertex redundantly rigid, because removing $k$ vertex results in removing at least $k$ edges. Therefore, there exists a vertex $v\in V$ where $G^{v}=G\setminus\{v\}$ is a flexible graph. Let U = {$u_{1},\cdots,u_{m}$} a set of all nodes where $u_{i}\in G^{v}$ and $e_{i}=(u_{i},v)\in E$. In addition, since $G$ is not $k$-edge ($k\geq 2$) redundantly rigid, we can select $v$ and $u_{m}$ so that subgraph $G^{e_{m}}=G\setminus\{(u_{m},v)\}$ is flexible. In summary, we consider a (fault satellite) $v$ and another vertex $e_{m}$ that satisfies
  • (A) $G^{v}=G\setminus\{v\}$ is flexible, and
  • (B) $G^{e_{m}}=G\setminus\{(u_{m},v)\}$ is flexible.
We prove that $G$ is embeddable by induction.
i) First, we prove $G^{v}\cup e_{1}$ is embeddable, almost surely
Proof: It is proved that for almost all realizations of the flexible graph, the flexing space of $f_{r}$ contains a submanifold that is diffeomorphic to the circle (Hendrickson,, 1992). Therefore, since $G^{v}$ is flexible from assumption (A), the distance between node $u_{1}$ and $v$ will be a multivariate function for almost every point on this circle. Hence, set of distances between $u_{1}$ and $v$ corresponding to $f_{r}$, denoted by $S_{d(u_{i},v)}(r)$ has measure non-zero. Thus, with probability greater than 0, $W(e_{1})\in S_{d(u_{1},v)}$, meaning $G^{v}\cup e_{1}$ is embeddable.
ii) Next, we prove that if $\tilde{G}^{i}=G^{v}\cup e_{1}\cup\cdots\cup e_{i}\ (1\leq i\leq m-1)$ is embeddable, $\tilde{G}^{i+1}=G^{v}\cup e_{1}\cup\ldots\cup e_{i}\cup e_{i+1}$ is embeddable, almost surely
Proof: $\tilde{G}^{i}$ is flexible since it is a subgraph of a flexible graph $G^{e_{m}}=G\setminus\{(u_{m},v)\}$ (from condition (B)). Therefore, with probability greater than 0, $W(e_{i+1})\in S_{d(u_{i+1},v)}$. Hence, given $\tilde{G}^{i}$ is embeddable, $\tilde{G}^{i+1}=\tilde{G}^{i}\cup{e_{i+1}}$ is also embeddable.
From i) and ii), $G=\tilde{G}^{m}$ is embeddable. Therefore (non k-vertex rigid graph) $G$ contains fault and is embeddable, meaning it is not fault disprovable. ∎
The visual description of the proof for $\mathbb{R}^{2}$ is shown in Figure 3.
Figure 3: A visual description of the proof of Proposition 2.3 for $d=2$. We cannot detect faults by looking at the embeddability of graphs that are not 2-edge redundantly rigid.
2.2.3 Summary
From proposition 2.2 and 2.3, we obtain the following proposition.
Proposition 2.4.
Given a weighted graph $G$, $G$ is fault disprovable if $G$ is k-vertex redundantly rigid ($k\geq 2)$, and only if $G$ is k-edge redundantly rigid ($k\geq 2)$.
@@ 3 Properties of the Euclidean Distance Matrix
In the previous section, we proved that by using vertex redundantly rigid graphs, we can detect faults by analyzing their embeddability. Analyzing the embeddability of an arbitrary graph is known to be NP-hard  (Hendrickson,, 1992), but if the graph is fully connected, we can analyze them by analyzing the ranks of the GCEDM. Fully connected graphs of more than 5 nodes are 2-vertex redundantly rigid in $\mathbb{R}^{3}$.
In this section, we introduce and prove several properties of the EDM and GCEDM constructed from the observed ranges between the satellites. We show how we can use the singular values of the GCEDM to evaluate if a graph is embeddable in $\mathbb{R}^{3}$.
3.1 Geometric Centered Euclidean Distance Matrix
Consider a fully connected graph of $n$ nodes. Using the set of observed ranges $r_{ij}$, we construct an EDM, where its elements are equivalent to the square of the observed ranges ($r_{ij}^{2})$. Let $\mathbf{D}^{n,d,m}\in\mathbb{R}^{n\times n}$ be an EDM without measurement noise ($w_{ij}=0$) and $m$ nodes out of $n$ nodes being fault. For example, a noiseless EDM with 6 satellites (in 3D space) with 2 fault satellites is denoted as $\mathbf{D}^{6,3,2}$. Similarly, we define $\tilde{\mathbf{D}}^{n,d,m}\in\mathbb{R}^{n\times n}$ as EDM with measurement noise and $m$ nodes out of $n$ nodes having a fault. A GCEDM $\mathbf{G}^{n,d,m}$ (or $\tilde{\mathbf{G}}^{n,d,m}$ for noisy EDM $\tilde{\mathbf{D}}^{n,d,m}$) is constructed from the EDM from the following operation